# PAtt-Lite: Lightweight Patch and Attention MobileNet for Challenging Facial Expression Recognition

### v1

@@ Abstract
Facial Expression Recognition (FER) is a machine learning problem that deals with recognizing human facial expressions. While existing work has achieved performance improvements in recent years, FER in the wild and under challenging conditions remains a challenge. In this paper, a lightweight patch and attention network based on MobileNetV1, referred to as PAtt-Lite, is proposed to improve FER performance under challenging conditions. A truncated ImageNet-pre-trained MobileNetV1 is utilized as the backbone feature extractor of the proposed method. In place of the truncated layers is a patch extraction block that is proposed for extracting significant local facial features to enhance the representation from MobileNetV1, especially under challenging conditions. An attention classifier is also proposed to improve the learning of these patched feature maps from the extremely lightweight feature extractor. The experimental results on public benchmark databases proved the effectiveness of the proposed method. PAtt-Lite achieved state-of-the-art results on CK+, RAF-DB, FER2013, FERPlus, and the challenging conditions subsets for RAF-DB and FERPlus. The source code for the proposed method will be available at https://github.com/JLREx/PAtt-Lite.
@@
Facial Expression Recognition, Patch Extraction, Self-Attention, MobileNetV1.
1 Introduction
@@
The main challenges that FER poses differently from other image classification tasks are the inter-class similarities and intra-class differences in human facial expressions. Inter-class similarities refer to the subtle differences between facial expressions, which makes it difficult to highlight the small differences between facial expressions and recognize them correctly. On the other hand, intra-class differences, also known as subject variability, refer to the characteristic of FER databases that images from an expression class are made up of different subjects with different facial structures, gender, age, and race. This variability can hinder the learning performance of a solution, as the model may struggle to generalize across different subjects, leading to reduced accuracy and reliability. For example, the differences between an angry face and a disgusted face may be minimal, while the differences between two different individuals within the same expression class can be quite significant.
@@
2 Related Work
2.1 Convolutional Neural Network
Convolutional Neural Network (CNN) is a class of deep learning models designed to process grid-like data such as images. It employs convolutional layers to automatically detect patterns or features through spatial hierarchies, enabling the network to learn progressively complex information, and pooling layers to reduce dimensionality and computational complexity, while maintaining important features. These networks have achieved significant results in computer vision tasks such as image classification, image segmentation, and object detection. The main advantage of CNNs is their ability to learn complex features automatically without the need for manual feature engineering. Besides, CNNs are also highly adaptable to different input sizes while being able to handle complex patterns and data variations.
LeNet-5[6] was the first CNN introduced in 1998 for the handwritten digit recognition problem, which has laid a foundation for modern CNN architecture. The CNN architecture has then gained its popularity with the introduction of AlexNet[7] in the 2012 ImageNet Large Scale Visual Recognition Challenge (ILSVRC). Since then, novel CNN architectures like the GoogLeNet[8], ResNet[9], Inception-v3[10], MobileNet[11], and more have been proposed for the general image classification task.
With the advancement in GPU technology and the availability of mature deep learning libraries, existing work for FER has focused more on deep learning solutions recently. These solutions often outperform the handcrafted methods, especially in in-the-wild databases. While many deep learning methods exist for FER, CNN has remained the more popular choice for existing work. Most of the CNN-based methods, such as [12, 13, 14, 5, 15, 16, 17], attempt to improve the FER performance by exploiting local information in different ways with their additional modules.
The development of CNN architectures has brought forward many innovations, including residual connections[9], bottleneck design[8, 9], batch normalization[18] and its alternatives[19, 20], depthwise separable convolutions[11], and many more. However, the architectures that integrate some of these innovations are complex and have a higher number of training parameters, which in turn require larger databases and longer training time. Hence, the ImageNet-pre-trained MobileNetV1[11] is selected as the baseline architecture for this research, due to its high performance despite its lightweight nature and relatively simpler architecture.
@@ 2.2 Vision Transformers
Although the patch extraction block is ultimately inspired by ViT, there exist some differences in terms of the implementation details. The first difference is the design and placement of the patch extraction block. The patch extraction mechanism in ViT is a single-layer convolution that is placed at the beginning of the architecture, while the patch extraction block in the proposed PAtt-Lite is a multi-layer convolution that is placed within the architecture. This placement allows the proposed method to fully utilize the pretrained weights of the backbone MobileNetV1, which were trained on ImageNet samples of size $224\times 224$. Secondly, ViT splits the input image of size $224\times 224$ into 196 non-overlapping patches of size $16\times 16$, while PAtt-Lite splits the output feature maps from MobileNetV1 of size $14\times 14$ into 4 non-overlapping patches of size $1\times 1$. The larger receptive regions of the proposed patch extraction block also help the proposed PAtt-Lite in extracting significant and high-level facial features.
@@
3 Methodology
Figure 1: Architecture of the proposed PAtt-Lite.
3.1 Overview
@@ 3.2 MobileNetV1
Depthwise separable convolution is a departure from the standard convolutional operation, as they split a standard convolution into two separate operations by performing depthwise convolutions followed by pointwise convolutions. Depthwise convolution is different from conventional convolution in that depthwise convolution applies a single convolutional filter for each input channel, whereas conventional convolution has filters that are as deep as its input. Meanwhile, pointwise convolution can be achieved using the standard convolutional operation by setting the kernel size to 1. Essentially, pointwise convolutions enable the mixing of input channels as conventional convolutions do.
As a lightweight CNN architecture, MobileNetV1 is selected as the base feature extractor in the proposed method as it provides easy finetuning of the pre-trained weights on our benchmark databases without overfitting the training samples. Mathematically, the feature extractive process from the truncated MobileNetV1 is formulated as follows:
@@
3.3 Patch Extraction
The key advantage of using a pre-trained CNN for transfer learning is that earlier layers have learned generic features of the training samples, such as the edges, while the later layers have learned specific features of training samples. In the context of PAtt-Lite, layers after the depthwise convolution of block 9 are skipped. The patch extraction block is added to better adapt to the FER databases than simply fine-tuning the final layers. This modification to the feature extractor also results in a shorter training period, as a higher learning rate can be used as opposed to when the pre-trained weights are being finetuned.
Our proposed patch extraction block consists of three different convolutional layers, the first two being depthwise separable convolutional layers while the last being a pointwise convolutional layer. Operating on feature maps from the MobileNetV1 that are padded to the dimension of $16\times 16$, the first separable convolutional layer is responsible for splitting the feature maps into four patches while learning higher-level features from its input. Subsequently, the second separable convolutional layer and the pointwise convolutional layer are responsible for learning the higher-level features from the patched feature maps, resulting in output with a dimension of $2\times 2$. Instead of the standard convolutional layer used in conventional CNNs, the depthwise separable convolutional layer is selected for PAtt-Lite. This design decision improves the classification performance of the proposed method on challenging subsets while reducing the number of model parameters.
3.4 Global Average Pooling
@@ 3.5 Attention Classifier
4 Experiments and Comparison
4.1 Databases
Both laboratory-controlled and in-the-wild databases are used to evaluate the proposed PAtt-Lite, namely CK+, RAF-DB, FER2013, and FERPlus.
@@ 4.1 Databases
4.2 Implementation Details
@@
4.3 Ablation Study
For the ablation study, the effectiveness of the patch extraction block and the attention classifier are evaluated by comparing them to the baseline performance of the finetuned MobileNetV1. Experiments are conducted on all benchmark databases for a proper evaluation on the effect of the proposed modules. Additionally, experimental results for the comparison between patch extraction and patch attention on in-the-wild databases are also included to justify the proposed patch extraction instead of conventional patch attention. The experimental results for this study are presented in Table I and Table II.
TABLE I: Ablation study for the proposed method on all benchmark databases. The best result is highlighted in bold.
@@ 4.3 Ablation Study
4.3.1 Effectiveness of proposed modules
As shown in Table I, the proposed patch extraction block is observed to have a slight decrease in performance compared to the MobileNetV1 baseline in in-the-wild databases. The performance drops in in-the-wild databases are mainly due to the difference in the number of trainable parameters between the original layers and the layers from the patch extraction block, which the newly initialized layers also resulted in the absence of pre-trained weights from the final layers. On the other hand, the MobileNetV1 baseline struggled to get a 100.00% mean accuracy on CK+ based on our experiments. However, with the small scale of CK+, this performance is immediately achievable with the addition of the patch extraction block, hence validating its effectiveness.
@@ 4.3.1 Effectiveness of proposed modules
4.3.2 Comparison between patch extraction and patch attention
TABLE II: Comparison between patch extraction and patch attention on in-the-wild databases. The best result is highlighted in bold.
@@ 4.3.2 Comparison between patch extraction and patch attention
As shown in Table II, the conventional patch attention mechanism is not bringing any performance gain when compared to the proposed patch extraction block. Instead, experimental results have demonstrated that the proposed patch extraction block performed better in the proposed method than the patch attention across all in-the-wild databases.
While the purpose to integrate a patch attention mechanism is often to highlight significant local facial regions and hence improve the classification performance under challenging conditions, the experimental results and confusion matrices depicted in Fig. 2 - 5 have shown that the patch attention is identical to the patch extraction in terms of performance under challenging conditions. Specifically, the classification accuracy and per-class performance under challenging conditions are similar between these two designs. Both performed similarly well overall and in most classes except the Contempt class and the Disgust class, where both struggled to perform well.
4.4 Method Comparison
In this section, PAtt-Lite is compared with state-of-the-art methods on the benchmark databases. The performance comparisons are shown in Table III - VII. In addition, the confusion matrices of the proposed PAtt-Lite on RAF-DB, FER2013, and FERPlus are depicted in Fig. 6.
@@ 4.4 Method Comparison
4.4.1 Results on CK+
TABLE III: Comparison of the state-of-the-art results on CK+. The best result is highlighted in bold.
@@ 4.4.1 Results on CK+
The performance comparison of the proposed method with the state-of-the-art methods for CK+ is presented in Table III. The proposed PAtt-Lite outperformed all CNN-based existing work[14, 15, 42, 43] and transformer-based existing work[26] in terms of cross-validation mean accuracy for CK+ by achieving 100.00% mean accuracy across 10-fold cross-validation. To the best of our knowledge, this is the best performance reported for a method that is trained without facial landmarks.
4.4.2 Results on RAF-DB.
TABLE IV: Comparison of the state-of-the-art results on RAF-DB and FERPlus. The best result is highlighted in bold.
@@
TABLE V: Per-class performance comparison of the state-of-the-art results on RAF-DB. The best result is highlighted in bold.
@@
The performance comparison of the proposed PAtt-Lite with state-of-the-art methods on RAF-DB is shown in Table IV. An accuracy and parameter comparison between the proposed method and state-of-the-art methods is also presented in Fig. 7. VTFF[31], TransFER[27], Facial Chirality[45], APViT[29], POSTER[28], and POSTER++[30] are transformer-based architecture, while RAN[5], SCAN-CCI[15], and ARM[44] are CNN-based architecture.
@@
Based on the comparison, the transformer-based methods are generally outperforming and have a greater number of parameters than the CNN-based methods, with SCAN-CCI[15] being the exception as it has 70M parameters despite having a CNN-based architecture. Our proposed PAtt-Lite achieved better performance with a CNN backbone than all transformer-based state-of-the-art. Specifically, PAtt-Lite has 2.84% over transformer-based POSTER++[30], which despite being the lightest transformer-based method, recorded the best result for RAF-DB. For comparison with CNN-based methods, PAtt-Lite achieved an improvement of 4.63% overall accuracy over ARM[44]. In terms of performance comparison with lightweight methods, our proposed method outperformed RAN[5] and ARM[44] by 8.15% and 4.63% respectively, while having 10 times lesser parameters.
@@
Furthermore, following [31, 28, 29, 30], a per-class performance comparison is added in Table V to compare the performance of PAtt-Lite on different classes of the database. For [15, 16, 32, 27, 44], which did not specifically report their per-class accuracy, their results from the confusion matrices are taken for comparison. Overall, a similar trend for the per-class performance is observed when PAtt-Lite is compared with existing work, where the per-class accuracy for the Disgust class and the Fear class is lacking behind the other classes. However, our method managed to slightly improve the per-class accuracy for the Disgust class over APViT[29] and perform on par with it for the Fear class while having 59 times fewer parameters. Overall, small improvements can be observed across all expression classes except class Anger and class Neutral, where significant improvement is recorded for the former and a small performance drop is noted for the latter. Noticeably, the proposed method managed to record a 6.79% improvement over all other state-of-the-art methods for class Anger, while recording a 1.87% performance drop over ARM[44]. Although the per-class accuracy for the Disgust class and the Fear class can be further improved, PAtt-Lite demonstrated consistently strong performance by managing to achieve about 93% per-class accuracy or greater in the remaining classes, resulting in an average accuracy of 90.38%. This corresponds to an improvement in the average accuracy of 4.02% over APViT[29], which previously reported the best average accuracy. Moreover, when compared to Imponderous Net[16], which has the nearest number of parameters as our proposed method, PAtt-Lite achieved an improvement of 12.81% in terms of average accuracy, while performing stronger and more consistent on all expression classes.
@@
4.4.3 Results on FER2013
TABLE VI: Comparison of the state-of-the-art results on FER2013. The best result is highlighted in bold.

### v2

@@ Abstract
Facial Expression Recognition (FER) is a machine learning problem that deals with recognizing human facial expressions. While existing work has achieved performance improvements in recent years, FER in the wild and under challenging conditions remains a challenge. In this paper, a lightweight patch and attention network based on MobileNetV1, referred to as PAtt-Lite, is proposed to improve FER performance under challenging conditions. A truncated ImageNet-pre-trained MobileNetV1 is utilized as the backbone feature extractor of the proposed method. In place of the truncated layers is a patch extraction block that is proposed for extracting significant local facial features to enhance the representation from MobileNetV1, especially under challenging conditions. An attention classifier is also proposed to improve the learning of these patched feature maps from the extremely lightweight feature extractor. The experimental results on public benchmark databases proved the effectiveness of the proposed method. PAtt-Lite achieved state-of-the-art results on CK+, RAF-DB, FER2013, FERPlus, and the challenging conditions subsets for RAF-DB and FERPlus.
@@
Facial Expression Recognition, MobileNetV1, Patch Extraction, Self-Attention.
[…]
=-21pt
I Introduction
@@
The main challenges that FER poses differently from other image classification tasks are the inter-class similarities and intra-class differences in human facial expressions. Inter-class similarities refer to the subtle differences between facial expressions, which makes it difficult to highlight the small differences between facial expressions and recognize them correctly. On the other hand, intra-class differences, also known as subject variability, refer to the characteristic of FER databases that images from an expression class are made up of different subjects with different facial structures, gender, age, and race. This variability can hinder the learning performance of a solution, as the model may struggle to generalize across different subjects, leading to reduced accuracy and reliability. For example, the differences between an angry face and a disgusted face may be minimal, whereas the differences between two different individuals within the same expression class can be quite significant.
@@
II Related Work
II-A Convolutional Neural Network
Convolutional Neural Network (CNN) is a class of deep learning models designed to process grid-like data such as images. It employs convolutional layers to automatically detect patterns or features through spatial hierarchies, enabling the network to learn progressively complex information, and pooling layers to reduce dimensionality and computational complexity while maintaining important features. These networks have achieved significant results in computer vision tasks such as image classification, image segmentation, and object detection. The main advantage of CNNs is their ability to learn complex features automatically without the need for manual feature engineering. Besides, CNNs are also highly adaptable to different input sizes while being able to handle complex patterns and data variations.
With the advancement in GPU technology and the availability of mature deep learning libraries, existing work for FER has focused more on deep learning solutions recently. These solutions often outperform the handcrafted methods, especially in in-the-wild databases. Most of the CNN-based methods, such as [6, 7, 8, 5, 9, 10, 11], attempt to improve the FER performance by exploiting local information in different ways with their additional modules.
The development of CNN architectures has brought forward many innovations, including residual connections[12], bottleneck design[13, 12], batch normalization[14] and its alternatives[15, 16], depthwise separable convolutions[17], and many more. However, the architectures that integrate some of these innovations are complex and have a higher number of training parameters, which in turn require larger databases and longer training time. By contrast, the architecture of MobileNetV1[17] is simpler and lighter than most of the well-known CNN architectures by comparison. Thus, the ImageNet-pre-trained MobileNetV1 is selected as the baseline architecture for this research, due to its high performance despite its lightweight and simple architecture. This simple architecture also has provided an easy finetuning process since overfitting and underfitting on the benchmark databases are easy to control with this architecture.
@@ II-B Vision Transformers
Although the patch extraction block is ultimately inspired by ViT, there exist some differences in terms of the implementation details. The first difference is the design and placement of the patch extraction block. The patch extraction mechanism in ViT is a single-layer convolution that is placed at the beginning of the architecture, whereas the patch extraction block in the proposed PAtt-Lite is a multi-layer convolution that is placed within the architecture. This placement allows the proposed method to fully utilize the pretrained weights of the backbone MobileNetV1, which were trained on ImageNet samples of size $224\times 224$. Secondly, ViT splits the input image of size $224\times 224$ into 196 non-overlapping patches of size $16\times 16$, whereas PAtt-Lite splits the output feature maps from MobileNetV1 of size $14\times 14$ into 4 non-overlapping patches of size $1\times 1$. The larger receptive regions of the proposed patch extraction block also help the proposed PAtt-Lite in extracting significant and high-level facial features.
@@
III Methodology
Figure 1: Architecture of the proposed PAtt-Lite. The image sample will first go through the truncated MobileNetV1 for feature extraction, in which the output feature maps will be padded and used as input for the proposed patch extraction block. The output feature maps of dimensions $2\times 2\times 256$ from the patch extraction block will then be global average pooled before being taken by the attention classifier.
III-A Overview
@@ III-B MobileNetV1
Depthwise separable convolution is a departure from the standard convolutional operation, as they split a standard convolution into two separate operations by performing depthwise convolutions followed by pointwise convolutions. Depthwise convolution is different from conventional convolution in that depthwise convolution applies a single convolutional filter for each input channel, whereas conventional convolution has filters that are as deep as its input. Meanwhile, pointwise convolution can be achieved using the standard convolutional operation by setting the kernel size to 1. Effectively, pointwise convolutions enable the mixing of input channels as conventional convolutions do.
The architecture of MobileNetV1 is simpler than most of the well-known CNN architectures by comparison. Other than being a lightweight architecture, the absence of complex designs like residual connections and bottleneck layers also contributed to the easiness of finetuning the pre-trained weights on our benchmark databases optimally. Mathematically, the feature extractive process from the truncated MobileNetV1 is formulated as follows:
@@
III-C Patch Extraction
The key advantage of using a pre-trained CNN for transfer learning is that earlier layers have learned generic features of the training samples, such as the edges, whereas the later layers have learned specific features of training samples. In the context of PAtt-Lite, layers after the depthwise convolution of block 9 are skipped. The patch extraction block is added to better adapt to the FER databases than simply fine-tuning the final layers. This modification to the feature extractor also results in a shorter training period, as a higher learning rate can be used as opposed to when the pre-trained weights are being finetuned.
Our proposed patch extraction block consists of three different convolutional layers, the first two being depthwise separable convolutional layers and the last being a pointwise convolutional layer. Operating on feature maps from the MobileNetV1 that are padded to the dimension of $16\times 16$, the first separable convolutional layer is responsible for splitting the feature maps into four patches while learning higher-level features from its input. Subsequently, the second separable convolutional layer and the pointwise convolutional layer are responsible for learning the higher-level features from the patched feature maps, resulting in output with a dimension of $2\times 2$. Instead of the standard convolutional layer used in conventional CNNs, the depthwise separable convolutional layer is selected for PAtt-Lite. This design decision improves the classification performance of the proposed method on challenging subsets while reducing the number of model parameters.
The design process of the patch extraction block started with a grid search for the optimum MobileNetV1 output layer with the number of patches for the patch extraction block. The baseline block consisted of a convolutional layer and a pointwise convolutional layer, which is retained in the final design. Our grid search has experimented with five convolutional kernel sizes, which are 3, 4, 5, 7, 8, and all layers of MobileNetV1 with the feature map size of $2\times 2\times D$. The summary of these experiments is included in Table V.
The baseline patch extraction block is then redesigned based on the optimum patch size and output layer. With the pointwise convolutional layer kept as is, the first convolutional layer was swapped out for a separable convolutional layer, which has a significantly smaller number of parameters while still providing the same spatial and channel convolution. Meanwhile, another separable convolutional layer was added to the patch extraction block to keep the number of patches constant while removing the need for a larger convolutional kernel. Thus, we arrived at our final design for the patch extraction block.
III-D Global Average Pooling
@@ III-E Attention Classifier
IV Experiments and Comparison
IV-A Databases
Both laboratory-controlled and in-the-wild databases are used to evaluate the proposed PAtt-Lite, namely CK+, RAF-DB, FER2013, and FERPlus. Summaries of the class data distribution for in-the-wild databases and their challenging subsets are presented in Table I and Table II. A summary of the training distribution for in-the-wild databases is also shown in Table III.
@@ IV-A Databases
TABLE I: Summary of class data distribution for in-the-wild databases.
Databases Anger Disgust Fear Happiness Neutral Sadness Surprise Contempt
RAF-DB 867 877 355 5957 3204 2460 1619 -
FER2013 4953 547 5121 8989 6198 6077 4002 -
FERPlus 3123 253 825 9367 13014 4414 4493 221
TABLE II: Summary of class data distribution for challenging subsets of RAF-DB and FERPlus.
Subsets Anger Disgust Fear Happiness Neutral Sadness Surprise Contempt
RAF-DB
Occlusion 31 47 35 236 191 122 72 -
Pose 30 75 78 38 446 282 179 149 -
Pose 45 37 36 20 164 129 90 82 -
FERPlus
Occlusion 23 3 33 122 162 125 135 2
Pose 30 102 5 28 286 466 151 127 5
Pose 45 52 1 16 141 271 89 59 4
TABLE III: Summary of training distribution for in-the-wild databases.
Databases Training Validation Testing
RAF-DB 12271 - 3068
FER2013 28659 3584 3582
FERPlus 28558 3579 3573
IV-B Implementation Details
@@
IV-C Ablation Study
For the ablation study, the effectiveness of the patch extraction block and the attention classifier are evaluated by comparing them to the baseline performance of the finetuned MobileNetV1. Experiments are conducted on all benchmark databases for a proper evaluation on the effect of the proposed modules. The summary of the experimental results for our grid search on the optimum output layer and number of patches is also included in Table V. Additionally, experimental results for the comparison between patch extraction and patch attention on in-the-wild databases are also included to justify the proposed patch extraction instead of conventional patch attention. The experimental results for this study are presented in Table IV and Table VI.
TABLE IV: Ablation study for the proposed method on all benchmark databases. The best result is highlighted in bold.
@@ IV-C Ablation Study
IV-C1 Effectiveness of proposed modules
As shown in Table IV, the proposed patch extraction block is observed to have a slight decrease in performance compared to the MobileNetV1 baseline in in-the-wild databases. The performance drops in in-the-wild databases are mainly due to the difference in the number of trainable parameters between the original layers and the layers from the patch extraction block, which the newly initialized layers also resulted in the absence of pre-trained weights from the final layers. On the other hand, the MobileNetV1 baseline struggled to get a 100.00% mean accuracy on CK+ based on our experiments. However, with the small scale of CK+, this performance is immediately achievable with the addition of the patch extraction block, hence validating its effectiveness.
@@ IV-C1 Effectiveness of proposed modules
IV-C2 Patch Extraction Block
Based on the experimental results in Table V, output feature maps from activated depthwise convolutional layers are the optimum layers for our patch extraction block, whereby patch sizes of 7 and 8 (with padding around the feature maps) generally yield the best results. Hence, this means splitting the feature maps into four patches is more optimal for our implementations.
TABLE V: Summary of classification accuracy for different output layers and different kernel sizes on RAF-DB. The best accuracy is highlighted in bold.
Layer Kernel Best Accuracy (%)
17 NP7 90.03
17 0P3 89.96
17 0P5 89.80
17 P4 89.11
17 P8 88.92
23 NP7 93.71
23 P8 93.45
23 0P5 91.69
23 0P3 91.17
23 P4 89.93
29 NP7 94.78
29 0P3 93.84
29 P8 93.64
29 P4 89.41
29 0P5 80.22
35 P8 94.56
35 NP7 93.64
35 P4 90.51
35 0P5 85.43
35 0P3 82.95
41 P8 94.39
41 NP7 93.09
41 0P5 92.86
41 P4 89.50
41 0P3 82.66
47 P8 87.84
47 NP7 86.38
47 0P5 84.03
47 0P3 79.04
47 P4 78.13
However, as the kernel size for the convolution operation gets bigger, so does its number of parameters, which contradicts with our idea of designing a small and lightweight model for facial expression recognition. Hence, we derived the design of the patch extraction block by experimenting with the replacement of the convolutional layer with a large kernel for the depthwise separable convolutional layer with a smaller kernel. In theory, this should still retain the performance of the original design as the separable convolutional layer is still performing the spatial and channel convolution that a conventional convolutional layer has. We also intend to keep the number of patches constant while trying to use a smaller kernel. Thus, the padding to the output feature maps from MobileNetV1 was kept, and a patch size of 4 was used for the first separable convolutional layer, and a patch size of 2 was used for the second separable convolutional layer. From our experimental results, the replacement proved to be successful with the tradeoff of a minor performance drop for a significant reduction in the number of parameters.
IV-C3 Comparison between patch extraction and patch attention
TABLE VI: Comparison between patch extraction and patch attention on in-the-wild databases. The best result is highlighted in bold.
@@ IV-C3 Comparison between patch extraction and patch attention
Figure 5: Confusion matrices of patch attention on challenging subsets of FERPlus.
As shown in Table VI, the conventional patch attention mechanism does not bring any performance gain when compared to the proposed patch extraction block. Instead, experimental results have demonstrated that the proposed patch extraction block performed better in the proposed method than the patch attention across all in-the-wild databases.
Whereas the purpose of integrating a patch attention mechanism is often to highlight significant local facial regions and hence improve the classification performance under challenging conditions, the confusion matrices depicted in Fig. 2 - 5 have shown that the patch attention is identical to the patch extraction in terms of performance under challenging conditions. Specifically, the classification accuracy and per-class performance under challenging conditions are similar between these two designs. Both performed similarly well overall and in most classes except the Contempt class and the Disgust class, where both struggled to perform well.
IV-D Method Analysis
Figure 6: Grad-CAM of PAtt-Lite on all seven classes of RAF-DB. The first row is the sample images from the testing set whereas the second row is the Grad-CAM visualizations for each sample from the row above.
Figure 7: Examples of true and false predictions on RAF-DB and their activation maps. The ground truths of the samples are labeled in black whereas the true and false predictions are labeled in green and red, respectively.
In this section, PAtt-Lite is analyzed using Grad-CAM visualizations in Fig. 6 and Fig. 7. It can be observed that PAtt-Lite can recognize the facial expression of the subjects through or around the occlusion and under a variety of pose angles, besides some that have weird ratios.
As highlighted in [39], we believe expression annotations are the limiting factor of today’s FER methods. Upon close inspection of the Grad-CAM from Fig. 7, it can be observed that some predictions, such as the fifth sample on the third row and the second sample on the fourth row, appear to be more closely represented by predictions from PAtt-Lite rather than the ground truths. These bad annotations are not limited to the test set but are available in the training set as well, which we believe contributed to PAtt-Lite’s overfitting to the databases, as can be observed in Fig. 7, where some samples, such as the sixth samples on the second row, may seem ambiguous or incorrectly predicted.
Besides the annotation, we also noticed that PAtt-Lite generally performed better on faces that are well-posed instead of faces that have a weird ratio, messy, or require context information. This is also illustrated in the visualizations in Fig. 7, such as the first sample on the third row and the fourth sample on the fourth row, where the backbone extractor randomly highlights the feature maps instead of focusing on the facial features.
On the other hand, the lower class accuracy in smaller classes can be explained by referring to Table I, Table II, and the confusion matrices, which most pointed that our method generally performed worse in the smaller classes. When the percentages are converted into the number of samples, it can be clearly observed that PAtt-Lite made a similar number of false predictions across all classes.
IV-E Method Comparison
In this section, PAtt-Lite is compared with state-of-the-art methods on the benchmark databases. The performance comparisons are shown in Table VII - XI. In addition, the confusion matrices of the proposed PAtt-Lite on RAF-DB, FER2013, and FERPlus are depicted in Fig. 8.
@@ IV-E Method Comparison
IV-E1 Results on CK+
TABLE VII: Comparison of the state-of-the-art results on CK+. The best result is highlighted in bold.
@@ IV-E1 Results on CK+
The performance comparison of the proposed method with the state-of-the-art methods for CK+ is presented in Table VII. The proposed PAtt-Lite outperformed all CNN-based existing work[8, 9, 40, 41] and transformer-based existing work[23] in terms of cross-validation mean accuracy for CK+ by achieving 100.00% mean accuracy across 10-fold cross-validation. To the best of our knowledge, FER-VT[42] is the only other method that reported the same performance with a transformer-based method.
IV-E2 Results on RAF-DB.
TABLE VIII: Comparison of the state-of-the-art results on RAF-DB and FERPlus. The best result is highlighted in bold.
@@
TABLE IX: Per-class performance comparison of the state-of-the-art results on RAF-DB. The best result is highlighted in bold.
@@
The performance comparison of the proposed PAtt-Lite with state-of-the-art methods on RAF-DB is shown in Table VIII. An accuracy and parameter comparison between the proposed method and state-of-the-art methods is also presented in Fig. 9. VTFF[28], TransFER[24], Facial Chirality[45], APViT[26], POSTER[25], POSTER++[27], and ARBEx[47] are transformer-based architecture, whereas RAN[5], SCAN-CCI[9], ARM[44], EAC[43], and DDAMFN[46] are CNN-based architecture.
@@
Based on the comparison, the transformer-based methods are generally outperforming and have a greater number of parameters than the CNN-based methods, with SCAN-CCI[9] being the exception as it has 70M parameters despite having a CNN-based architecture. Our proposed PAtt-Lite achieved better performance with a CNN backbone than all transformer-based state-of-the-art. Specifically, PAtt-Lite has 2.58% over transformer-based ARBEx[47], a modified version of the lightest transformer-based method, POSTER++[27], recorded the best result for RAF-DB. For comparison with CNN-based methods, PAtt-Lite achieved an improvement of 4.63% overall accuracy over ARM[44]. In terms of performance comparison with lightweight methods, our proposed method also outperformed RAN[5], EAC[43], ARM[44], and DDAMFN[46] by 8.15%, 5.06%, 4.63%, and 3.70% respectively, while having significantly lesser parameters.
@@
Furthermore, following [28, 25, 26, 27], a per-class performance comparison is added in Table IX to compare the performance of PAtt-Lite on different classes of the database. For [9, 10, 29, 24, 44], which did not specifically report their per-class accuracy, their results from the confusion matrices are taken for comparison. Overall, a similar trend for the per-class performance is observed when PAtt-Lite is compared with existing work, where the per-class accuracy for the Disgust class and the Fear class is lacking behind the other classes. However, our method managed to slightly improve the per-class accuracy for the Disgust class over APViT[26] and perform on par with it for the Fear class while having 59 times fewer parameters. Overall, small improvements can be observed across all expression classes except class Anger and class Neutral, where significant improvement is recorded for the former and a small performance drop is noted for the latter. Noticeably, the proposed method managed to record a 6.79% improvement over all other state-of-the-art methods for class Anger, while recording a 1.87% performance drop over ARM[44]. Although the per-class accuracy for the Disgust class and the Fear class can be further improved, PAtt-Lite demonstrated consistently strong performance by managing to achieve about 93% per-class accuracy or greater in the remaining classes, resulting in an average accuracy of 90.38%. This corresponds to an improvement in the average accuracy of 4.02% over APViT[26], which previously reported the best average accuracy. Moreover, when compared to Imponderous Net[10], which has the nearest number of parameters as our proposed method, PAtt-Lite achieved an improvement of 12.81% in terms of average accuracy, while performing stronger and more consistent on all expression classes.
@@
IV-E3 Results on FER2013
TABLE X: Comparison of the state-of-the-art results on FER2013. The best result is highlighted in bold.